OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6.

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

\[y-(-1)=6(x-5)\] \[y+1=6(x-5)\]is the point slope form